20. A patient is to receive 600 units of heparin per hour. The drop factor is 60 microdrops per milliliter. The order states that 12,500 units of heparin are to be added to 250 milliliters of 5% D/W, a.How many microdrops per minute will you infuse? 21. A patient has an IV of 1000 milliliters of 5% D/W The flow rate is 36 drops per minute. The drop factor is 10 drops per milliliter a.How many hours will it take for this IV to finish? 22. An infusion of 1000 milliliters of lactated Ringers solution is infusing at a rate of 17 drops per minute What time will it be completed if the drop factor is 15 drops per milliliter and this infusion started at 7 am. a. 23. How many hours will an IV of 1000 milliliters of 5% D/W infuse if the order is for 125 milliliters per hour? 24. The thrombolytic enzyme streptokinase (Streptase) has been ordered intravenously, 1,500,000 units over 60 minutes in 50 milliliters of 0.9% NS. a. What will the flow rate be in microdrops per minute? 25. The prescriber has ordered 250 milliliters of 5% D/W with 250 milligrams of esmolol hydrochloride (Brevbod at a rate of 10 micrograms per kilogram per minute IV What will the flow rate be in milliliters per hour if the patient weighs 75 kilograms? 26. If an IV is infusing at a rate of 80 microdrops per minute, how many milliliters per hour will you patient receive? 27. The prescriber has ordered 500 milliliters of 5% D/W with 500 milligrams of Lidocaine IV and your patient, who weighs 80 kilograms, must receive 20 micrograms per kilogram per minute. What will the flow rate be in militers per hour? a.
Answer #1

20.volume to be infused =dose required /dose available*volume of solution

=600/12500*250=12ml

Drops/min=volumeto be infused /time in minutes*drop factor=12ml/60min*60=12microdrops/min

21.time required to finish the infusion=Total volume to be infused /flow rate *drop factor=1000/36*10=277.78min=4.6hours

22.volume to be infused =1000ml

Rate=17 drops /min

Drop factor=15drops/ml

Duration=total volume (ml)/rate*dropfactor

=1000/17*15=882min=14 hrs 42 min

Starting time of infusion=7am

Finishing time=7am +14 hrs 42 min=9.42pm

23.total volume =1000ml

Rate =125ml/hr

Total time =volume /rate=1000/125=8hours